solve each system. substitute a for 1/x and b for 1/y. solve for a and b, then find x and y
1/x+2/y=-1
2/x-1/y=-7

Answer by uday1435(57)

1/x+2/y=-1
2/x-1/y=-7
Substitutig a = 1/x and b =1/y we get
a + 2b = - 1 Eq(1)
2a – b = - 7 Eq(2) .
Multiply Eq(1) by 2 and subtract from Eq.(2) we get
2a – b = - 7 Eq(2) .
2a + 4b = - 2 Multiplied Eq(1) by 2
- 5b = - 5 subtractig the above eqs
b = (- 5/- 5) = 1
Pluging in the value of b in Eq(1) we have
a + 2(1) = -1 ; a = - 1 – (2) = - 3; so x =1/a = 1/(-3) = - 1/3 and y = 1/b = 1/1 =1
Hence the solution is a = - 1/3 and b = 1